# From Features to Actions: Explainability in Traditional and Agentic AI Systems

### v1

@@ Abstract
Over the last decade, explainable AI has primarily focused on interpreting individual model predictions, producing post-hoc explanations that relate inputs to outputs under a fixed decision structure. Recent advances in large language models (LLMs) have enabled agentic AI systems whose behaviour unfolds over multi-step trajectories. In these settings, success and failure are determined by sequences of decisions rather than a single output. While useful, it remains unclear how explanation approaches designed for static predictions translate to agentic settings where behaviour emerges over time. In this work, we bridge the gap between static and agentic explainability by comparing attribution-based explanations with trace-based diagnostics across both settings. To make this distinction explicit, we empirically compare attribution-based explanations used in static classification tasks with trace-based diagnostics used in agentic benchmarks (TAU-bench Airline and AssistantBench). Our results show that while attribution methods achieve stable feature rankings in static settings (Spearman ρ = 0.86), they cannot be applied reliably to diagnose execution-level failures in agentic trajectories. In contrast, trace-grounded rubric evaluation for agentic settings consistently localizes behaviour breakdowns and reveals that state tracking inconsistency is 2.7× more prevalent in failed runs and reduces success probability by 49%. These findings motivate a shift towards trajectory-level explainability for agentic systems for evaluating and diagnosing autonomous AI behaviour.
@@ 1 Introduction
However, the AI landscape has fundamentally changed. The rise of agentic AI systems [19], particularly Large Language Model (LLM)-based agents, represents a departure from the static, single-inference settings for which most traditional explainability methods were developed. Historically, Explainable AI methods (XAI) has focused on providing post-hoc explanations for individual predictions, typically under the assumption of a fixed input-output mapping and a single decision point, while interpretability has referred more broadly to models or representations that are transparent by design [32]. In contrast, agentic systems operate through sequences of observations, decisions, and tool invocations that unfold over dozens of steps before producing a final outcome [13]. This shift from prediction to action, and from single inference to extended reasoning, challenges core assumptions of traditional XAI , rendering them inadequate for explaining agentic systems’ behaviour.
Problem Statement. In static settings, explanations are commonly framed as attributions over input features or tokens for a single prediction. In agentic settings, many critical questions are trajectory-level: why an agent chose a tool, abandoned a strategy, propagated an incorrect state, or failed to recover after an error. In other words, the unit of explanation is no longer a single prediction but a trajectory, a sequence of states, actions, and observations that collectively determines success or failure. We argue that bridging static and agentic explainability requires re-conceptualizing explanation targets from feature-level influence to trajectory-level decision accounts, and coupling explanation artifacts with execution context and faithfulness signals grounded in the run.
@@ 3 Explainability Across Static and Agentic AI Paradigms
Explainability methods have evolved alongside advances in AI, from early rule-based systems and decision trees to post-hoc feature attributions, concept-based analyses, and mechanistic interpretability for deep neural networks. The rise of tool-using, multi-step agentic systems introduces new explainability requirements: explanations must account for action sequences, tool interactions, and state evolution over time. In this section, we formalize the distinction between static and agentic paradigms, map existing XAI families through this lens, and introduce a unified framework—the Minimal Explanation Packet (MEP)—for packaging explanation artifacts with context and verification when behaviour unfolds over trajectories. Figure 1 conceptually contrasts static and agentic MEPs, highlighting the shift from single input–output explanations to trajectory-level accounts.
@@ 3.2 Mapping of XAI Methods Across Paradigms
CoT- and ReAct-style traces expose intermediate reasoning and make multi-step structure legible to humans. However, they are self-reported and can diverge from the actual causal drivers of behaviour, especially when tool results, memory updates, or retrieval events dominate subsequent decisions. Trace-based explanations therefore become more reliable when aligned with interaction logs—linking each stated intent to the corresponding action, tool output, and state update—and when discrepancies between “said” and “done” are surfaced explicitly.
@@ 4.1 Datasets
LLM judge GPT-5 Medium (single-pass, trace-only)
@@
Inference models GPT-4.1 (AssistantBench), o4-mini-2025-04-16 (TAU-bench)
@@ 4.2 Models
In the static setting, we use two lightweight classifiers: TF–IDF with Logistic Regression and a Text CNN baseline. In the agentic setting, we study tool-using LLM agents on TAU-bench Airline and AssistantBench, using o4-mini-2025-04-16 [30] and GPT-4.1 as the underlying models, respectively. All agent traces are labeled post-hoc with Docent [27], using a fixed GPT-5 Medium judge for consistent rubric evaluation.
@@ 5.2 Agentic Explainability
We analyze agent behaviour using execution traces collected with HAL-Harness on two benchmarks: TAU-bench Airline and AssistantBench. Our goal is to diagnose why agents succeed or fail beyond aggregate accuracy. Each trace is labeled using the behavioural rubrics defined in Section 3.3, with a fixed GPT-5 Medium judge  [^2] https://openai.com/index/gpt-5-system-card/ to ensure consistency.
@@ 5.2.2 Trace-Level Rubric Analysis
In trace-level rubric analysis, we label each execution trace with behavioural rubrics provided in Section 3.3. To ensure consistency, we use a single fixed judge (GPT-5 Medium [29]) in a single pass conditioned only on traces. The results on trace-level rubric analysis are discussed next.
@@ 5.3 Bridging Static and Agentic Explainability
Each trajectory is first labelled using Docent rubrics (Section 3.3). and encoded as a compact binary feature vector, where each dimension indicates whether a behavioural constraint is satisfied or violated. Using these rubric features, we train a logistic regression model to predict task success vs. failure (hyperparameters in Table 4). We compute SHAP values using a linear explainer consistent with the model to quantify the influence of each rubric feature on the surrogate outcome predictor.

### v2

@@ Abstract
Over the last decade, XAI has primarily focused on interpreting individual model predictions, producing post-hoc explanations that relate inputs to outputs under a fixed decision structure. Recent advances in large language models (LLMs) have enabled agentic AI systems whose behaviour unfolds over multi-step trajectories. In these settings, success and failure are determined by sequences of decisions rather than a single output. While useful, it remains unclear how explanation approaches designed for static predictions translate to agentic settings where behaviour emerges over time. In this work, we bridge the gap between static and agentic explainability by comparing attribution-based explanations with trace-based diagnostics across both settings. To make this distinction explicit, we empirically compare attribution-based explanations used in static classification tasks with trace-based diagnostics used in agentic benchmarks (TAU-bench Airline and AssistantBench). Our results show that while attribution methods achieve stable feature rankings in static settings (Spearman ρ = 0.86), they cannot be applied reliably to diagnose execution-level failures in agentic trajectories. In contrast, trace-grounded rubric evaluation for agentic settings consistently localizes behaviour breakdowns and reveals that state tracking inconsistency is 2.7× more prevalent in failed runs and reduces success probability by 49%. These findings motivate a shift towards trajectory-level explainability for agentic systems for evaluating and diagnosing autonomous AI behaviour.
@@ 1 Introduction
However, the AI landscape has fundamentally changed. The rise of agentic AI systems [19], particularly LLM-based agents, represents a departure from the static, single-inference settings for which most traditional explainability methods were developed. Historically, Explainable AI methods (XAI) have focused on providing post-hoc explanations for individual predictions, typically under the assumption of a fixed input-output mapping and a single decision point, while interpretability has referred more broadly to models or representations that are transparent by design [33]. In contrast, agentic systems operate through sequences of observations, decisions, and tool invocations that unfold over dozens of steps before producing a final outcome [13]. This shift from prediction to action, and from single inference to extended reasoning, challenges core assumptions of traditional XAI , rendering them inadequate for explaining agentic systems’ behaviour.
Problem Statement. In static settings, explanations are commonly framed as attributions over input features or tokens for a single prediction. In agentic settings, many critical questions are trajectory-level: why an agent chose a tool, abandoned a strategy, propagated an incorrect state, or failed to recover after an error. In other words, the unit of explanation is no longer a single prediction but a trajectory: a sequence of states, actions, and observations that collectively determines success or failure. We argue that bridging static and agentic explainability requires re-conceptualizing explanation targets from feature-level influence to trajectory-level decision accounts, and coupling explanation artifacts with execution context and faithfulness signals grounded in the run.
@@ 3 Explainability Across Static and Agentic AI Paradigms
Explainability methods have evolved alongside advances in AI, from early rule-based systems and decision trees to post-hoc feature attributions, concept-based analyses, and mechanistic interpretability for deep neural networks. The rise of tool-using, multi-step agentic systems introduces new explainability requirements: explanations must account for action sequences, tool interactions, and state evolution over time. In this section, we formalize the distinction between static and agentic paradigms, map existing XAI families through this lens, and introduce a unified framework, the Minimal Explanation Packet (MEP), for packaging explanation artifacts with context and verification when behaviour unfolds over trajectories. Figure 1 conceptually contrasts static and agentic MEPs, highlighting the shift from single input–output explanations to trajectory-level accounts.
@@ 3.2 Mapping of XAI Methods Across Paradigms
CoT- and ReAct-style traces expose intermediate reasoning and make multi-step structure legible to humans. However, they are self-reported and can diverge from the actual causal drivers of behaviour, especially when tool results, memory updates, or retrieval events dominate subsequent decisions. Trace-based explanations therefore become more reliable when aligned with interaction logs, linking each stated intent to the corresponding action, tool output, and state update, and when discrepancies between “said” and “done” are surfaced explicitly.
@@ 4.1 Datasets
LLM judge GPT-5 (medium reasoning effort, single-pass, trace-only)
@@
Inference models GPT-4.1 (AssistantBench), o4-mini-2025-04-16 (TAU-bench Airline)
@@ 4.2 Models
In the static setting, we use two lightweight classifiers: TF–IDF with Logistic Regression and a Text CNN baseline. In the agentic setting, we study tool-using LLM agents on TAU-bench Airline and AssistantBench, using o4-mini-2025-04-16 [30] and GPT-4.1 [29] as the underlying models, respectively. All agent traces are labeled post-hoc with Docent [27], using a fixed GPT-5 [31] judge with medium reasoning effect for consistent rubric evaluation.
@@ 5.2 Agentic Explainability
We analyze agent behaviour using execution traces collected with HAL-Harness on two benchmarks: TAU-bench Airline and AssistantBench. Our goal is to diagnose why agents succeed or fail beyond aggregate accuracy. Each trace is labeled using the behavioural rubrics defined in Section 3.3, with a fixed GPT-5 judge  [^2] https://openai.com/index/gpt-5-system-card/ to ensure consistency.
@@ 5.2.2 Trace-Level Rubric Analysis
In trace-level rubric analysis, we label each execution trace with behavioural rubrics provided in Section 3.3. To ensure consistency, we use a single fixed judge (GPT-5 [31]) in a single pass conditioned only on traces. The results on trace-level rubric analysis are discussed next.
@@ 5.3 Bridging Static and Agentic Explainability
Each trajectory is first labelled using Docent rubrics (Section 3.3) and then encoded as a compact binary feature vector, where each dimension indicates whether a behavioural constraint is satisfied or violated. Using these rubric features, we train a logistic regression model to predict task success vs. failure (hyperparameters in Table 4). We compute SHAP values using a linear explainer consistent with the model to quantify the influence of each rubric feature on the surrogate outcome predictor.